OpenStudy (anonymous):

Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5... Answer -1/5 + 2/5 n 1/5 + 2/5 n 1 2/5 - 2/5 n

OpenStudy (anonymous):

Each successive term gets bigger by 2/5, right?

OpenStudy (anonymous):

And the sequence starts with 1/5

OpenStudy (anonymous):

1/5 +2/5n

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

grow a chest u manorexic guy

hartnn (hartnn):

its -1/5 + 2/5 n !!

OpenStudy (anonymous):

@hartnn, why would you start with -1/5 ? I don't follow...

hartnn (hartnn):

an=a+(n-1)d a=1st term=1/5 d=common difference=2/5 an=1/5 + (n-1)*(2/5) = 1/5 +2/5 n - 2/5 = -1/5 + 2n/5

OpenStudy (anonymous):

Oh... it's because the expression I ended up with gives the incorrect answer for term n=1. I was thinking about terms after 1/5 being the sequence , but the "normal" way (i.e. "correct" ;) ) is to call the first term the n=1 term. You're right... good catch!

OpenStudy (anonymous):

@dnance Check your sequence question again... @hartnn has the correct answer... I messed it up earlier... sorry!!